OpenStudy (anonymous):

If complex number z=4-6i,then |z-2|= ?

OpenStudy (shamim):

\[z=4-6i\]

OpenStudy (shamim):

\[z-2=4-6i-2=2-6i\]

OpenStudy (shamim):

\[|z-2|=\sqrt{2^{2}+(-6)^{2}}\]

OpenStudy (shamim):

now try it

OpenStudy (anonymous):

Im confused... Can you llease explain the process?

OpenStudy (shamim):

where u confused

OpenStudy (anonymous):

Where you took a square root of 2-6i

OpenStudy (shamim):

its rule of mode

OpenStudy (shamim):

its law of modulus

OpenStudy (shamim):

i took modulus on left side n root on right side

OpenStudy (anonymous):

Oh i thought the sticks were absolute value signs... Thanks anyway

OpenStudy (shamim):

welcome

OpenStudy (anonymous):

it is the same